Abstract

We investigate the coefficients generated by expressing the falling factorial (xy)\underline{k} as a linear combination of falling factorial products x\underline{l}y\underline{m} for l, m = 1, …, k. Algebraic and combinatoric properties are discussed, some in relation to Stirling numbers.